OpenStudy (anonymous):

solve the equation by completing the square 2x^2 = 3x + 1

OpenStudy (tkhunny):

2x^2 - 3x = 1 2(x^2 - (3/2)x + _______) = 1 + 2(________) (3/2)/2 = 3/4 (3/4)^2 = 9/16 2(x^2 - (3/2)x + (9/16)) = 1 + 2(9/16) Almost done.